# On the wall-normal velocity variance in canonical wall-bounded turbulence

### v1

@@ Abstract
The variance and spectra of wall-normal velocities are investigated for direct numerical simulations of turbulent flow in a channel, pipe, and zero-pressure-gradient (ZPG) boundary layer, across a decade of wall friction Reynolds numbers. Spectra along the spanwise wavenumber have a pronounced peak at intermediate wavenumbers that is proportional to the turbulent kinetic energy dissipation rate and the cube of the local shear stress throughout the bottom half of the boundary layer for all flow cases. The observed scaling with the local stress rather than the surface shear velocity $U_{\tau}$ accounts for differences in the ZPG wall-normal variance seen in previous studies. The scaling is attributed to the fact that wall-normal motions are predominately ‘active’ per Townsend’s attached eddy hypothesis and directly contribute to the local shear stress, while noting this hypothesis does not account for the linear decay of the total stress in enclosed flows. The Reynolds number dependence of the variance is determined from the scale separation between the spectrum peak and the dissipative cutoff, where the resulting semi-empirical fit aligns with variance values in the literature. At the high-Reynolds-number limit, the spectral peak leads to a variance between 1.45 to 1.65 times the local shear stress. This range is consistent with previous predictions relative to $U_{\tau}$, including for the vertical velocity in the near-neutral atmospheric boundary layer. However, universality in the exact proportional constant is precluded by relatively minor contributions from the ‘inactive’ motions at low wavenumbers, which vary with wall-normal position and for different flow configurations.
@@ 1 Introduction
The Attached Eddy Hypothesis (AEH) of A. A. Townsend (1976), along with several decades of subsequent investigations (Marusic & Monty, 2019), is a key advancement in describing turbulent statistics in canonical wall-bounded flows. The AEH provides a phenomenological link between the turbulence, the logarithmic mean velocity profile, and the logarithmic decay in the streamwise and spanwise velocity variances. These trends have all been supported by measurements at sufficiently high Reynolds numbers across various canonical wall-bounded flows (see, e.g., Jiménez & Hoyas, 2008; Hultmark et al., 2012; Marusic et al., 2013; Orlandi et al., 2015; Lee & Moser, 2015) including the near-neutral atmospheric surface layer (Puccioni et al., 2023; Qin et al., 2025).
The population of attached eddies envisioned by Townsend also produces a constant wall-normal velocity variance $\overline{w^{\prime 2}}$ and Reynolds shear stress $\overline{u^{\prime}w^{\prime}}$ within the logarithmic region of the boundary layer:
@@
where both are proportional to the surface shear (friction) velocity $U_{\tau}=\sqrt{\tau_{w}/\rho}$ corresponding to the wall shear stress $\tau_{w}$ and fluid density $\rho$. Wall-normal profiles of $\overline{w^{\prime 2}}$ tend to vary only weakly within the logarithmic region (Morrill-Winter et al., 2015; Qin et al., 2025) and there is evidence for a constant third-order skewness (Buono et al., 2024b). Yet, there is no conclusive universal value for $B_{3}$ from observations and Eq. (1) has the least experimental support among the statistics predicted by the AEH (Buono et al., 2024a).
@@
There are two primary challenges in determining the value of $B_{3}$. The first is that the AEH models an asymptotically high-Reynolds-number wall-bounded flow with negligible viscosity effects within the logarithmic region. Observations of $\overline{w^{\prime 2}}$ in available simulations and laboratory-scale experiments are limited to finite Reynolds number and must account for its dependence. An example is shown in Fig. 1(a) for direct numerical simulations (DNS) of flows in a zero-pressure-gradient (ZPG) flat plate boundary layer (Sillero et al., 2013), channel (Lee & Moser, 2015), and pipe (Yao et al., 2023). The $\overline{w^{\prime 2}}$ for channel and pipe flows is nearly constant across intermediate distances, but the magnitude increases with the friction Reynolds number $Re_{\tau}=U_{\tau}\delta/\nu$ with $\delta$ being the boundary layer thickness and $\nu$ the kinematic viscosity.
@@
Perry and co-authors (1986; 1990) estimated this Reynolds-number dependence by considering the finite scale separation between the wall parameters and the viscous motions in the spectrum of $w$, resulting in a correction for $B_{3}$ as a function of $z^{+}=zU_{\tau}/\nu$. An alternative approach was recently proposed by Spalart & Abe (2021), who used extrapolation of profiles at different Reynolds numbers following the technique of Luchini (2017). The method accurately describes channel data (Hoyas et al., 2022), but is empirical in nature and is not linked to the flow phenomenology. These previous efforts collectively have not produced a conclusive Reynolds number correction to Eq. (1) and further research on the matter is warranted.
@@
Figure 1: Profiles of the wall-normal velocity variance $\overline{{w^{\prime}}^{2}}$ normalized by the surface shear velocity $U_{\tau}$: (a) wall-normal position $z$ in viscous units $z^{+}=zU_{\tau}/\nu$; (b) $z$ in outer units relative to the boundary layer thickness $\delta$. In all figures, the data are from direct numerical simulations (DNS) (Sillero et al., 2013; Lee & Moser, 2015; Yao et al., 2023), with color indicating the flow case and shade corresponding to the friction Reynolds numbers $Re_{\tau}=\delta U_{\tau}/\nu$ given in Table 1.
@@
The exclusion of viscous effects in the AEH and Eq. (1) applies also to the Reynolds shear stress in Eq. (2), which neglects the viscous stress contribution. Equation (2) further neglects any decrease in the total stress profile below $U_{\tau}^{2}$ as the wall-normal distance increases. Given that $\overline{w^{\prime 2}}$ and $\overline{u^{\prime}w^{\prime}}$ have both been postulated to receive contributions solely from the ‘active’ motions per Townsend’s (1976) AEH, it is possible that viscous stresses and a decreasing total stress can directly impact $\overline{w^{\prime 2}}$. However, these effects have not been closely investigated in the context of Eq. (1).
@@
The second challenge to determining $B_{3}$ is an observed difference in the variance between wall-bounded flow configurations. As seen in Fig. 1, the ZPG boundary layer has higher $\overline{w^{\prime 2}}$ relative to $U_{\tau}$ and a peak that occurs farther from the surface. This difference is often discussed in terms of enclosed versus non-enclosed flows. Jiménez & Hoyas (2008) conducted a survey of available datasets and acknowledged the different amplitudes (see, e.g., their Fig. 4). A similar survey by Buschmann & Gad-el Hak (2009) discussed the different position of the variance peak, but noted limited differences in the amplitudes. An explanation for these trends has not been offered in these previous studies, however, and it remains an open question regarding the wall-normal variance.
@@
The combination of the Reynolds-number dependence and the different profile for ZPG boundary layers has led to a range of reported $B_{3}$ values for the high-$Re$ limit. Spalart (1988) found $B_{3}\approx$ 1.75 empirically from early DNS at low Reynolds numbers. Perry & Li (1990) refined their spectral approach with a model spectrum (Kovasznay, 1948) and fitted $B_{3}\approx 1.6$ to experimental ZPG boundary layer data. Hafez (1991) revised this value to $B_{3}\approx 1.78$ using further measurements, and the higher $B_{3}$ agrees well with measurements in high-$Re$ laboratory facilities (Morrill-Winter et al., 2015) as well as the atmospheric surface layer (Kunkel & Marusic, 2006). Empirical fits to a wider range of atmospheric measurements suggest $B_{3}\approx$ 1.55 to 1.7 in neutral conditions (Panofsky et al., 1977; Garratt, 1992). Recent experimental pipe flow data are also consistent with a higher $B_{3}$ value close to 1.85 (Örlü et al., 2017).
@@
The same DNS featured in the Fig. 1 example are used throughout the study. The focus of the analysis is to identify scaling behavior in the spectrum of the wall-normal velocity. The integrated spectrum is then used to assess the implications of the observed scaling on the variance statistics, including for the questions given above. The remainder of the article is outlined as follows: the methodology is detailed in §2, including an overview of the simulations and spectrum analysis; results are then presented in §3; the implications for the motivating questions are discussed in §4; concluding remarks are given in §5.
@@ 2.1 Simulations
Three published DNS datasets are used here: ZPG flat plate boundary layer (Sillero et al., 2013), channel flow (Lee & Moser, 2015), and pipe flow (Yao et al., 2023). Each dataset includes a range of friction Reynolds numbers that are listed in Table 1 and shown as different color shades in Fig. 1. All statistics and variables directly follow the published datasets, except where noted below.
@@
For the ZPG flat plate dataset, three intermediate runs between $Re_{\tau}$ = 1310 and 1990 are excluded here for clarity in later figures, and on account of the relatively modest $Re_{\tau}$ difference across all runs compared to the other cases. The reported boundary layer thickness $\delta_{99}$ based on the convergence of the mean flow to free-stream conditions is not used. Rather, $\delta$ is determined as the $z$ position where the local total shear stress is $\tau(z)=0.0015\tau_{w}$ based on $\tau_{w}$. Defining $\delta$ based on the vanishing stress is more appropriate for comparison between flow cases due to the relevance of the stress profile to the later analysis; for the enclosed cases, the stress is zero at the channel half-height and pipe center that correspond to $\delta$. The present TBL thickness definition yields $\delta$ values within 1% of estimates from the phenomenological definition given recently by Lozier et al. (2025), which is nominally $\gtrsim$ 25% than the value of $\delta_{99}$ (Morrill-Winter et al., 2015; Lozier et al., 2025).
@@ 2.2 Spanwise wavenumber spectra
One-dimensional velocity spectra are traditionally studied as a function of the streamwise wavenumber $k_{x}$. Along this direction – and also for time series – the wall-normal velocity spectrum $E_{ww}(k_{x})$ is approximately constant at the lowest wavenumbers before undergoing an extended transition to the inertial subrange of scales at higher $k_{x}$ (Ayet & Katul, 2020). The number of contributing factors to the shape of this transition make it infeasible to identify the appropriate length scale for this region of the spectrum.
In contrast, the spectrum $E_{ww}(k_{y})$ as a function of the spanwise wavenumber $k_{y}$ exhibits a distinct peak at intermediate scales as seen in Fig. 2(a). There is a clearer segregation of motions in the spanwise spectrum, where the peak and low-wavenumber limit have different behaviors and can be individually assessed. Accordingly, $E_{ww}(k_{y})$ statistics are investigated in this study. The spectra for the ZPG flat plate case are available at a limited number of $z$ positions, such that wall-normal profiles of spectral properties are coarse for this case in later figures.
Figure 2: Example wall-normal velocity spectra $E_{ww}$ as a function of spanwise wavenumber $k_{y}$ for $z/\delta=$ 0.1, where the dashed lines are fits using the model von Kármán spectrum in Eq. (3). (a) for the $Re_{\tau}=$ 5 200 channel flow case (Lee & Moser, 2015) to demonstrate the wavenumber $k_{peak}$ and energy density $E_{peak}$ of the spectrum peak, the low-wavenumber plateau $E_{k_{y}\to 0}$, and the contribution of large-scale motions $\overline{{w^{\prime}}^{2}}_{l}=\int_{0}^{k_{peak}}E_{ww}dk_{y}$ and small-scale motions $\overline{{w^{\prime}}^{2}}_{s}=\int_{k_{peak}}^{\infty}E_{ww}dk_{y}$ to the wall-normal variance. (b) for all cases in Table 1 to demonstrate the model spectrum fit.
@@
The $E_{ww}(k_{y})$ spectra are well described by a model von Kármán spectrum (Diederich & Drischler, 1957):
@@
where $a$, $b$, and $c$ are fitted. The dissipative cutoff occurring at the Kolmogorov microscale $\eta$ is $f(k_{y}\eta)=\exp(-5.2[\{(k_{y}\eta)^{4}+0.15^{4}\}^{1/4}-0.15])$ (Pope, 2000) with $\beta=$ 5.2 following experimental evidence (Saddoughi & Veeravalli, 1994), and 0.15 chosen empirically based on the alignment with the DNS spectra seen in Fig. 2(b).
Equation (3) is a general form of the von Kármán spectrum because the integral length $L$ and root-mean-square velocity are not imposed. The scaling properties of the spectrum are a goal of the analysis and no parameters are assumed aside from $\eta$. The primary use of the model in Eq. (3) is to estimate the low-wavenumber limit $E_{k_{y}\to 0}=a$ that corresponds to the energy splashing region captured within the DNS domain (Ayet & Katul, 2020; Katul et al., 2013). It is noted that the wavenumber $k_{peak}$ associated with the spectral peak may be determined analytically when assuming $f(k_{peak}\eta)=1$ (i.e. high Reynolds number flow). This wavenumber location is determined by setting $dE_{ww}(k_{y})/dk_{y}=0$ to yield
@@
This finding implies that (i) $b>11/6$, which is satisfied by $b=8/3$ in the original spectrum model (Diederich & Drischler, 1957); and (ii) $k_{peak}$ is insensitive to $a$. While the leading factor does not determine the wavenumber of the peak, $E_{ww}(k_{y}=0)=a=2\overline{{w^{\prime}}^{2}}L_{w}/\pi$ relates to both the variance and the integral length scale $L_{w}$ of $w$ defined from the autocorrelation function. This definition for $E_{ww}(k_{y}=0)$ and $a$ results directly from the spectrum’s relation to the autocorrelation, which corresponds to $L_{w}$ for $k_{y}=0$.
2.3 Scaling parameters
A direct outcome of the attached eddy population in the AEH is that $z$ and $U_{\tau}$ are the appropriate scales for the energy-containing production region in the spectrum of turbulence. These scales are accordingly used in previous studies on the spectrum of $w$ (Perry et al., 1986; Perry & Li, 1990), and are included in this analysis as a basis for comparison.
The scales $z$ and $U_{\tau}$ are strictly applicable to the logarithmic region in the high-Reynolds-number limit, and the parameters must be generalized outside of these conditions. To this end, Davidson and coauthors (2006; 2014) proposed a length scale $U_{\tau}^{3}/\epsilon$ based on the local dissipation rate $\epsilon$. This scale provides a mathematically consistent transition into the inertial subrange of the spectrum where $\epsilon$ is the sole scaling parameter, and accounts for the fact that lengths besides $z$ may contribute to setting $\epsilon$ depending on the conditions (Davidson & Krogstad, 2014). In practice, $U_{\tau}^{3}/\epsilon$ asymptotes to $\kappa z$ in the high-$Re$ logarithmic region (de Silva et al., 2015), where $\kappa$ is the von Kármán constant, but outperforms $z$ in scaling the start of the inertial sublayer for finite Reynolds numbers (Davidson et al., 2006; Heisel et al., 2022), within the roughness sublayer (Davidson & Krogstad, 2014; Ghannam et al., 2018), and for stratified flow (Chamecki et al., 2017).
The velocity scale can also be generalized by considering a local-in-$z$ shear velocity corresponding to the decaying stress profile. This is a relaxation of Eq. (2) in the AEH to account for conditions when the local stress decreases below the surface value given by $U_{\tau}$. The generalization allows $\overline{w^{\prime 2}}$ to maintain its relation with the local stress as discussed in §1. In atmospheric flows, this approach is known as “local scaling” and it provides improved mean velocity similarity under stratified conditions (Nieuwstadt, 1984; Holtslag & Nieuwstadt, 1986; Heisel & Chamecki, 2023). The local velocity scale is used here also to define the dissipation-based length scale discussed above, leading to the following revised scaling parameters:

### v2

@@ Abstract
The variance and spectra of wall-normal velocities are investigated for direct numerical simulations of turbulent flow in a channel, pipe, and zero-pressure-gradient boundary layer across a decade of friction Reynolds numbers. Spectra along the spanwise wavenumber have a pronounced peak well described by the turbulent dissipation rate and the local shear stress throughout the bottom half of the boundary layer. Deviations in the local stress from the surface shear velocity $U_{\tau}$ account for almost all of the differences in wall-normal velocity variance observed across different canonical flows, including for plane Couette flow. The dependence on the local stress is attributed to the fact that wall-normal motions are predominately ‘active’ per Townsend’s attached eddy hypothesis and directly contribute to the local shear stress, noting this hypothesis assumes simplified ideal conditions with constant turbulent shear stress. A semi-empirical fit applied to the Reynolds number dependence of the variance matches the simulations across the lower half of the boundary layer and aligns with observed values in the literature. The fit extrapolates to a value between 1.45 and 1.65 times the local shear stress in the high-Reynolds-number limit, consistent with previous predictions relative to $U_{\tau}$ including for the vertical velocity in the near-neutral atmospheric boundary layer. However, universality in the exact proportional constant is precluded by small discrepancies in the variances corresponding to dissimilarity in the low-wavenumber contributions across different flow configurations and wall-normal positions. We speculate the dissimilarity is due to relatively weak ‘inactive’ wall-normal motions that are excluded from Townsend’s original hypothesis.
@@ 1 Introduction
The Attached Eddy Hypothesis (AEH) of A. A. Townsend (1976), along with several decades of subsequent investigations (Marusic and Monty, 2019), is a key advancement in describing turbulent statistics in canonical wall-bounded flows. The AEH provides a phenomenological link between the turbulence, the logarithmic mean velocity profile, and the logarithmic decay in the streamwise and spanwise velocity variances. These trends have all been supported by measurements at sufficiently high Reynolds numbers across various canonical wall-bounded flows (see, e.g., Jiménez and Hoyas, 2008; Hultmark et al., 2012; Marusic et al., 2013; Orlandi et al., 2015; Lee and Moser, 2015) including the near-neutral atmospheric surface layer (Puccioni et al., 2023; Qin et al., 2025).
The population of attached eddies envisioned by Townsend also produces a constant wall-normal velocity variance $\overline{w^{\prime 2}}$ within the logarithmic region of the boundary layer:
@@
which is proportional to the surface shear (friction) velocity $U_{\tau}=\sqrt{\tau_{w}/\rho}$ corresponding to the wall shear stress $\tau_{w}$ and fluid density $\rho$. Wall-normal profiles of $\overline{w^{\prime 2}}$ tend to vary only weakly within the logarithmic region (Morrill-Winter et al., 2015; Qin et al., 2025) and there is evidence for a constant third-order skewness (Buono et al., 2024b). Yet, there is no conclusive universal value for $B_{3}$ from observations and Eq. (1) has the least experimental support among the statistics predicted by the AEH (Buono et al., 2024a).
@@
There are two primary challenges in determining the value of $B_{3}$. The first is that the AEH models an asymptotically high-Reynolds-number wall-bounded flow with negligible viscosity effects within the logarithmic region. Observations of $\overline{w^{\prime 2}}$ in available simulations and laboratory-scale experiments are limited to finite Reynolds number and must account for its dependence. An example is shown in Fig. 1(a) for direct numerical simulations (DNS) of flows in a zero-pressure-gradient (ZPG) flat plate boundary layer (Sillero et al., 2013), channel (Lee and Moser, 2015), and pipe (Yao et al., 2023). The $\overline{w^{\prime 2}}$ for channel and pipe flows is nearly constant across intermediate distances, but the magnitude increases with the friction Reynolds number $Re_{\tau}=U_{\tau}\delta/\nu$ with $\delta$ being the boundary layer thickness and $\nu$ the kinematic viscosity.
@@
Perry and co-authors (1986; 1990) estimated this Reynolds-number dependence by considering the finite scale separation between the wall parameters and the viscous motions in the spectrum of $w$, resulting in a correction for $B_{3}$ as a function of $z^{+}=zU_{\tau}/\nu$. An alternative approach was recently proposed by Spalart and Abe (2021), who used extrapolation of profiles at different Reynolds numbers following the technique of Luchini (2017). The method accurately describes channel data (Hoyas et al., 2022), but is empirical in nature and is not linked to the flow phenomenology. These previous efforts collectively have not produced a conclusive Reynolds number correction to Eq. (1) and further research on the matter is warranted.
@@
Figure 1: Profiles of the wall-normal velocity variance $\overline{{w^{\prime}}^{2}}$ normalized by the surface shear velocity $U_{\tau}$: (a) wall-normal position $z$ in viscous units $z^{+}=zU_{\tau}/\nu$; (b) $z$ in outer units relative to the boundary layer thickness $\delta$. In all figures, the data are from direct numerical simulations (DNS) (Sillero et al., 2013; Lee and Moser, 2015; Yao et al., 2023), with color indicating the flow case and shade corresponding to the friction Reynolds numbers $Re_{\tau}=\delta U_{\tau}/\nu$ given in Table 1.
The exclusion of viscous effects in the AEH applies also to the Reynolds shear stress $\overline{u^{\prime}w^{\prime}}$. The assumption of a constant turbulent stress with negligible viscous stress contribution is an ansatz of the AEH:
$\dfrac{-\overline{u^{\prime}w^{\prime}}(z)}{U_{\tau}^{2}}=1.$ (2)
@@
Equation (2) further neglects any decrease in the total stress profile below $U_{\tau}^{2}$ as the wall-normal distance increases. Given that $\overline{w^{\prime 2}}$ and $\overline{u^{\prime}w^{\prime}}$ have both been postulated to receive contributions solely from the ‘active’ motions per Townsend’s (1976) AEH, viscous stresses and a decreasing total stress can directly impact $\overline{w^{\prime 2}}$. This has been demonstrated using simulations of channel flows under different Reynolds numbers and forcing conditions, where variability in $\overline{w^{\prime 2}}$ throughout most of the boundary layer follows $\overline{u^{\prime}w^{\prime}}(z)$ more closely than $U_{\tau}$ (Tuerke and Jiménez, 2013; Lozano-Durán and Bae, 2019). However, the effect of the stress profile has not been closely investigated in the context of $B_{3}$ and Eq. (1).
@@
The second challenge to determining $B_{3}$ is an observed difference in the variance between wall-bounded flow configurations. As seen in Fig. 1, the ZPG boundary layer has higher $\overline{w^{\prime 2}}$ relative to $U_{\tau}$ and a peak that occurs farther from the surface. This difference is often discussed in terms of enclosed versus non-enclosed flows. Jiménez and Hoyas (2008) conducted a survey of available datasets and acknowledged the different amplitudes (see, e.g., their Fig. 4). A similar survey by Buschmann and Gad-el-Hak (2009) discussed the different wall-normal position of the variance peak, but noted limited differences in the amplitudes. An explanation for these trends has not been offered in these previous studies, however, and it remains an open question regarding the wall-normal variance.
@@
The combination of the Reynolds-number dependence and the different profile for ZPG boundary layers has led to a range of reported $B_{3}$ values for the high-$Re$ limit. Spalart (1988) found $B_{3}\approx$ 1.75 empirically from early DNS at low Reynolds numbers. Perry and Li (1990) refined their spectral approach with a model spectrum (Kovasznay, 1948) and fitted $B_{3}\approx 1.6$ to experimental ZPG boundary layer data. Hafez (1991) revised this value to $B_{3}\approx 1.78$ using further measurements, and the higher $B_{3}$ agrees well with measurements in high-$Re$ laboratory facilities (Morrill-Winter et al., 2015) as well as the atmospheric surface layer (Kunkel and Marusic, 2006). Empirical fits to a wider range of atmospheric measurements suggest $B_{3}\approx$ 1.55 to 1.7 in neutral conditions (Panofsky et al., 1977; Garratt, 1992). Recent experimental pipe flow data are also consistent with a higher $B_{3}$ value close to 1.85 (Örlü et al., 2017).
@@
The same DNS featured in the Fig. 1 example are used throughout the study. The focus of the analysis is to identify the dependencies of the spectrum for the wall-normal velocity. The integrated spectrum is then used to assess the implications of the observed dependencies on the variance statistics, including for the questions given above. The remainder of the article is outlined as follows: the methodology is detailed in §2, including an overview of the simulations and spectrum analysis; results are then presented in §3; the implications for the motivating questions are discussed in §4; concluding remarks are given in §5.
@@ 2.1 Simulations
Three published DNS datasets are used here: ZPG flat plate boundary layer (Sillero et al., 2013), channel flow (Lee and Moser, 2015), and pipe flow (Yao et al., 2023). Each dataset includes a range of friction Reynolds numbers that are listed in Table 1 and shown as different color shades in Fig. 1. All statistics and variables directly follow the published datasets, except where noted below.
@@
For the ZPG flat plate dataset, three intermediate runs between $Re_{\tau}$ = 1310 and 1990 are excluded here for clarity in later figures, and on account of the relatively modest $Re_{\tau}$ difference across all runs compared to the other cases. The reported boundary layer thickness $\delta_{99}$ based on the convergence of the mean flow to free-stream conditions is not used. Rather, $\delta$ is determined as the $z$ position where the local total shear stress is $\tau(z)=0.0015\tau_{w}$. Defining $\delta$ based on the vanishing stress is more appropriate for comparison between flow cases due to the relevance of the stress profile to the later analysis; for the enclosed cases, the stress is zero at the channel half-height and pipe center that correspond to $\delta$. The present TBL thickness definition yields $\delta$ values within 1% of estimates from the phenomenological definition given recently by Lozier et al. (2025), which is nominally $\gtrsim$ 25% than the value of $\delta_{99}$ (Morrill-Winter et al., 2015; Lozier et al., 2025).
@@ 2.2 Spanwise wavenumber spectra
One-dimensional velocity spectra are traditionally studied as a function of the streamwise wavenumber $k_{x}$. Along this direction – and also for time series – the wall-normal velocity spectrum $E_{ww}(k_{x})$ is approximately constant at the lowest wavenumbers before undergoing an extended transition to the inertial subrange of scales at higher $k_{x}$ (Ayet and Katul, 2020). The number of contributing factors to the shape of this transition make it infeasible to identify the appropriate length scale for this region of the spectrum.
In contrast, the spectrum $E_{ww}(k_{y})$ as a function of the spanwise wavenumber $k_{y}$ exhibits a distinct peak at intermediate scales as seen in Fig. 2(a). There is a clearer segregation of motions in the spanwise spectrum, where the peak and low-wavenumber limit have different behaviors and can be individually assessed. Accordingly, $E_{ww}(k_{y})$ statistics are investigated in this study. The later analysis was repeated using the peak in the premultiplied spectrum $k_{y}E_{ww}$ which is commonly employed in the literature for streamwise wavenumber spectra that otherwise lack a peak (see, e.g., Baidya et al., 2017). The same spectral scaling is observed for both approaches, but the uncompensated spectrum is less affected by the viscous cutoff and is the more robust option for this study. The spectra for the ZPG flat plate case are available at a limited number of $z$ positions, such that wall-normal profiles of spectral properties are coarse for this case in later figures.
Figure 2: Example wall-normal velocity spectra $E_{ww}$ as a function of spanwise wavenumber $k_{y}$ for $z/\delta=$ 0.1, where the dashed lines are fits using the model von Kármán spectrum in Eq. (3). (a) for the $Re_{\tau}=$ 5 200 channel flow case (Lee and Moser, 2015) to demonstrate the wavenumber $k_{peak}$ and energy density $E_{peak}$ of the spectrum peak, the low-wavenumber plateau $E_{k_{y}\to 0}$, and the contribution of large-scale motions $\overline{{w^{\prime}}^{2}}_{l}=\int_{0}^{k_{peak}}E_{ww}dk_{y}$ and small-scale motions $\overline{{w^{\prime}}^{2}}_{s}=\int_{k_{peak}}^{\infty}E_{ww}dk_{y}$ to the wall-normal variance. (b) for all cases in Table 1 to demonstrate the model spectrum fit.
@@
The $E_{ww}(k_{y})$ spectra are well described by a model von Kármán spectrum (Diederich and Drischler, 1957):
@@
where $a$, $b$, and $c$ are fitted. The dissipative cutoff occurring at the Kolmogorov microscale $\eta$ is $f(k_{y}\eta)=\exp(-\beta[\{(k_{y}\eta)^{4}+0.15^{4}\}^{1/4}-0.15])$ (Pope, 2000) with $\beta=$ 5.2 following experimental evidence (Saddoughi and Veeravalli, 1994), and 0.15 chosen empirically based on the alignment with the DNS spectra seen in Fig. 2(b).
Equation (3) is a general form of the von Kármán spectrum because the integral length $L$ and root-mean-square velocity are not imposed. The scaling properties of the spectrum are a goal of the analysis and no parameters are assumed aside from $\eta$. The primary use of the model in Eq. (3) is to estimate the low-wavenumber limit $E_{k_{y}\to 0}=a$ that corresponds to the energy splashing region captured within the DNS domain (Ayet and Katul, 2020; Katul et al., 2013). It is noted that the wavenumber $k_{peak}$ associated with the spectral peak may be determined analytically when assuming $f(k_{peak}\eta)=1$ (i.e. high Reynolds number flow). This wavenumber location is determined by setting $dE_{ww}(k_{y})/dk_{y}=0$ to yield
@@
This finding implies that (i) $b>11/6$, which is satisfied by $b=8/3$ in the original spectrum model (Diederich and Drischler, 1957); and (ii) $k_{peak}$ is insensitive to $a$. While the leading factor does not determine the wavenumber of the peak, $E_{ww}(k_{y}=0)=a=2\overline{{w^{\prime}}^{2}}L_{w}/\pi$ relates to both the variance and the integral length scale $L_{w}$ of $w$ defined from the autocorrelation function. This definition for $E_{ww}(k_{y}=0)$ and $a$ results directly from the spectrum’s relation to the autocorrelation, which corresponds to $L_{w}$ for $k_{y}=0$.
2.3 Flow parameters
A direct outcome of the attached eddy population in the AEH is that $z$ and $U_{\tau}$ are the appropriate scales for the energy-containing production region in the spectrum of turbulence. These scales are accordingly used in previous studies on the spectrum of $w$ (Perry et al., 1986; Perry and Li, 1990), and are included in this analysis as a basis for comparison.
The scales $z$ and $U_{\tau}$ are strictly applicable to the logarithmic region in the high-Reynolds-number limit, and the parameters must be generalized outside of these conditions. To this end, Davidson and coauthors (2006; 2014) proposed a length $U_{\tau}^{3}/\epsilon$ based on the local dissipation rate $\epsilon$. This length provides a mathematically consistent transition into the inertial subrange of the spectrum where $\epsilon$ is the sole scaling parameter, and accounts for the fact that lengths besides $z$ may contribute to setting $\epsilon$ depending on the conditions (Davidson and Krogstad, 2014). In practice, $U_{\tau}^{3}/\epsilon$ asymptotes to $\kappa z$ in the high-$Re$ logarithmic region (de Silva et al., 2015), where $\kappa$ is the von Kármán constant, and in these conditions the premultiplied peak in the $w$ spectrum is well described by $z$ (Kunkel and Marusic, 2006; Baidya et al., 2017). However, $U_{\tau}^{3}/\epsilon$ outperforms $z$ in delineating the start of the inertial sublayer for lower Reynolds numbers (Davidson et al., 2006; Heisel et al., 2022), within the roughness sublayer (Davidson and Krogstad, 2014; Ghannam et al., 2018), and for stratified flow (Chamecki et al., 2017).
The velocity scale can also be generalized by considering a local-in-$z$ shear velocity corresponding to the decaying stress profile (Tuerke and Jiménez, 2013; Lozano-Durán and Bae, 2019). This is a relaxation of Eq. (2) in the AEH to account for conditions when the local stress decreases below the surface value given by $U_{\tau}$. The generalization allows $\overline{w^{\prime 2}}$ to maintain its relation with the local stress as discussed in §1. In atmospheric flows, this approach is known as “local scaling” and it provides improved mean velocity similarity under stratified conditions (Nieuwstadt, 1984; Holtslag and Nieuwstadt, 1986; Heisel and Chamecki, 2023). The local shear velocity is used here also to define the dissipation-based length scale discussed above, leading to the following revised scaling parameters: